OpenStudy (anonymous):

The equation for a projectile's height versus time is h(t)=-16t^2+v^0t+h^o. A tennis ball machine serves a ball vertically into the air from a height of 2 feet, with an initial speed of 120 feet per second. What is the maximum height, in feet, the ball will attain? Round to the nearest whole foot.

OpenStudy (jdoe0001):

\(\bf \text{initial velocity}\\ h = -16t^2+v_ot+h_o \qquad \text{in feet}\\ v_o = \textit{initial speed}\\ h_o = \textit{initial height}\\ \quad \\ \qquad thus\\ h = -16t^2+v_ot+h_o \implies h = -16t^2+(120)t+(2)\ \quad \\\implies h = -16t^2+120t+2\) notice you have a quadratic equation, and the graph is that one of a parabola, the leading coefficient is negative, thus the parabola is opening downwards, 

OpenStudy (jdoe0001):

you can find the vertex of a parabola equation like so, at the coordinates of \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\) you're asked to find the maximum height, thus the "y" value at the vertex only, so, you'd only need the y-coordinate

OpenStudy (jdoe0001):

a, b, and c, that is \(\bf y = ax^2+bx+c\)

OpenStudy (anonymous):

really dont get any of this.... so whats the answer after putting it all together?

OpenStudy (jdoe0001):

the maximum height, is "y-coordinate" at the vertex that is the highest point of the parabola